09) I now have $100 and three choices of investment during the next four years. For every dollar invested: Investment A yields $0.10 a year after the investment and yie years after the investment, Investment B yields $0.20 a year S1.10 two years after the investment, Investment C yields $1.50 th Investments A and C are available in the beginning of year I and 2 while B is available in the lds $1.30 three after the investment and yields ree years after the investment. During each year, uninvested casl funds, which yield 6% interest per year. At most $50 may be placed in each of investments A, B, and C in each year. Formulate an LP to maximize my cash on hand four years from now.
Answer #1

Let,

Bi = beginning cash at year-i beginning
Ei = ending cash at year-i end
Ai = investment in A in year-i beginning
Bi = investment in B in year-i beginning
Ci = investment in C in year-i beginning
Mi = investment in money market in year-i beginning
Si = total investment at the year-i beginning
Ri = total income at the year-i end

The situation is best depicted in the following tabular form.

Year
1 2 3 4
Beginning cash (Bi) B1 = 100 B2 = E1 B3 = E2 B4 = E3
Invest in,
A A1 A2 A3 A4
B B1 B2 B3 B4
C C1 C2 C3 C4
Money market M1 M2 M3 M4
Total invested (Si) S1 S2 S3 S4
Income from,
A 0.1A1 0.1A2 0.1A3 + 1.3A1 0.1A4 + 1.3A2
B 0.2B1 0.2B1 + 1.1B1 0.2B3 + 1.1B2 0.2B4 + 1.1B3
C 0 0 1.5C1 1.5C2
Money market 0.06M1 0.06M2 0.06M3 0.06M4
Total income (Ri) R1 R2 R3 R4
Ending cash (Ei) E1=B1+R1-S1 E2=B2+R2-S2 E3=B3+R3-S3 E4=B3+R4-S4

LP formulation

Maximize. Z = E4

Subject to,

B1 = 100
Bi = Ei-1 for i=2,3,4
Si = Ai + Bi + Ci + Mi for i=1,2,3,4
Ei = Bi + Ri - Si for i=1,2,3,4

R1 = 0.1A1 + 0.2B1 + 0.06M1
R2 = 0.1A2 + 0.2B1 + 1.1B1 + 0.06M2
R3 = 0.1A3 + 1.3A1 + 0.2B3 + 1.1B2 + 1.5C1 + 0.06M3
R4 = 0.1A4 + 1.3A2 + 0.2B4 + 1.1B3 + 1.5C2 + 0.06M4

Bi ≥ Si

Ai ≤ 50; Bi ≤ 50; Ci ≤ 50

Bi, Ei, Si, Ai, Bi, Ci, Mi, Ri ≥ 0